whats the answer to 6(k+2)+2(2k-5)=22
Answer by checkley77(12844)

6(k+2)+2(2k-5)=22
6k+12+4k-10=22
10k=22+10-12
10k=20
k=20/10
k=2 ans.
Proof:
6(2+2)+2(2*2-5)=22
6*4+2(4-5)=22
24+2*-1=22
24-2=22
22=22